OpenStudy (anonymous):

How does meiosis I differ from meiosis II? A. The homologous chromosomes pair up during meiosis II but not during meiosis I. B. Two sex cells are produced as a result of meiosis II but not as a result of meiosis I. C. The sister chromatids separate during meiosis I but not during meiosis II. D. Chromosome number decreases by half as a result of meiosis I but not as a result of meiosis II.

OpenStudy (abhisar):

I think its D

OpenStudy (abhisar):

what do u say @deepika.comet ?

OpenStudy (deepika.comet):

gave you a medal.... so i say you are right....:)

OpenStudy (abhisar):

oh yes..!!

OpenStudy (deepika.comet):

why cant it be C

OpenStudy (abhisar):

What do u say @Somy ?

OpenStudy (abhisar):

do sister chromatid get separated in meosis I @deepika.comet ?

OpenStudy (deepika.comet):

yeah no so actual reduction takes place in anaphase II im having a doubt in this when will the exact reduction divisiontakes place...? @Abhisar

OpenStudy (somy):

D i think

OpenStudy (abhisar):

i think separation of sister chromatids take place in continued mitosis.

OpenStudy (abhisar):

Or Meiosis II @Somy

OpenStudy (deepika.comet):

see exhange of chromosomes takes place in meiosis I [prophase I] but when will the exact reduction takes place resulting in hapliod cells

OpenStudy (somy):

sister chromatids separate in both meiosis 1 and 2

OpenStudy (deepika.comet):

then will will reduction of chromosomes takes place @somy

OpenStudy (somy):

yes it will

OpenStudy (abhisar):

no she meant when will

OpenStudy (abhisar):

It will take place in M I

OpenStudy (abhisar):

m i correct ?

OpenStudy (somy):

yes but it will reduce twice actually saying

OpenStudy (somy):

wait let me check

OpenStudy (abhisar):

Yes it will be in M I

OpenStudy (abhisar):

reduction division takes place in M I, in M II its only homotypic division which does not effect the chromosome number

OpenStudy (deepika.comet):

yes @abhisar is correct M I shows reduction of chromosomes and the MII is similar to mitotis division

OpenStudy (deepika.comet):

then the M I gives hapliod from diploid and M II gives hapliod only [ resembling mitosis]

OpenStudy (somy):

father + mother 23 + 23 = 46 (crossing over) 46/2= 23 each of the 2 cells produced then in these two cells 23*2= 46 then 46/2= 23 further 2 cells

OpenStudy (somy):

so basically saying reduction division happens twice in both 1 and 2 meiosis

OpenStudy (deepika.comet):

does autosomes participate in meiosis..?

OpenStudy (somy):

this is basically whats happening with number of chromosomes i don't really know which one to consider as 'reduction'

OpenStudy (deepika.comet):

ok reduction never occurs in both stages if so it results in chromosomes with reduced hapliods also..

OpenStudy (deepika.comet):

why did you multiply with 2 there

OpenStudy (somy):

because this happens the make a copy of genetic info which is from parents and then divide again

OpenStudy (somy):

http://www.youtube.com/watch?v=rqPMp0U0HOA i don't know what im saying already so just watch this

OpenStudy (abhisar):

Thankyou so much @Somy & @deepika.comet for ur help and support \(\huge\ddot\smile\)